OpenStudy (anonymous):

find an equation of the line describe below through (9,6) and (6,9) write the equation in slope-intercept solve for y when possible

OpenStudy (anonymous):

(9-6)/(6-9) 3/-3 = -1 = m y-y1=m(x-x1) y-6=-1(x-9) y-6=-x+9 y=-x+15

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

youre welcome